OpenStudy (chycora):

Which of the following expressions will produce an answer with three significant figures? Check all that apply. A. 100 • 3.4 B. 22.67 + 14.6 C. 20 - 40.8 D. 2101 / 23.7

OpenStudy (rcp031):

D. 2101 / 23.7

OpenStudy (chycora):

okay thanks

OpenStudy (chycora):

is that the only one

OpenStudy (rcp031):

significant figures: the figures of a number that express a magnitude to a specified degree of accuracy, rounding up or down the final figure: For example: 3.14159 to four significant figures is 3.142

OpenStudy (chycora):

so would it be B and D @rcp031

OpenStudy (rcp031):

A. 100 *3.4 = 340 mean three significant figure B. 22.67 + 14.6 mean 37.27 four significant C. 20 - 40.8=19.2 mean three significant figure D. 2101 / 23.7= mean 88.64879... so many significant figures

OpenStudy (rcp031):

My bad, actually correct answer is A and C, sorry for my mistake.

OpenStudy (chycora):

no it's okay thanks

OpenStudy (chycora):

it is B and D